OpenStudy (anonymous):

Factor the polynomial :5k^3+100k^2+500k

OpenStudy (anonymous):

factor out a k :P

OpenStudy (anonymous):

ooh better yet 5k since 100 and 500 are both divisible by 5k :D

OpenStudy (anonymous):

From the start, you can factor out 5k \[5 k (k^2+20 k+100)\]

OpenStudy (anonymous):

@timo86m okay and I get 5k(k^2 +20k+100) after that what do I do?

OpenStudy (anonymous):

you can further factor (k^2+20 k+100)

OpenStudy (anonymous):

Try to find two numbers that not only equal 20 when added but also equal 100 when multiplied together....

OpenStudy (anonymous):

Like 4 and 3: \[4 + 3 = 7\]\[4 \times 3 = 12\]

OpenStudy (anonymous):

its 10 and 10 =D

OpenStudy (anonymous):

try factroing a 10 since 10*10=100 10+10=20 (k+10)^2 is a factor of the above and you plug it back in to get 5 k (10+k)^2

OpenStudy (anonymous):

oh i was right!

OpenStudy (anonymous):

WOO-HOO! You got it

OpenStudy (anonymous):

and after that what's the next step?

OpenStudy (anonymous):

5 k (10+k)^2 Is factored :P

OpenStudy (anonymous):

Its completely factored now :)